# Two-Sample Testing with Block-Wise Missingness in Multi-Source Data

### v1

@@ 3.1 Finite-sample Properties
In this section, we provide explicit solution/formula to $\boldsymbol{\Sigma}_{v}$ and $\boldsymbol{\Sigma}_{c}$ under the null hypothesis, which is crucial to construct our proposed test statistics.
First, we define a collection of notations that are related to the rank matrix. Let $\delta_{\alpha\beta}=\mathbbm{1}(\alpha=\beta)$ be the Kronecker delta with the indicator function $\mathbbm{1}(\cdot)$. We further define:
$\bar{R}_{i\cdot}^{(\alpha\beta)}$ $=\frac{1}{N_{\beta}-\delta_{\alpha\beta}}\sum_{j=1}^{N_{\beta}}R_{ij}^{(\alpha\beta)},$ $r_{0}^{(\alpha\beta)}$ $=\frac{1}{N_{\alpha}}\sum_{i=1}^{N_{\alpha}}\bar{R}_{i\cdot}^{(\alpha\beta)},$
$(r_{1}^{(\alpha\beta\gamma)})^{2}$ $=\frac{1}{N_{\alpha}}\sum_{i=1}^{N_{\alpha}}\bar{R}_{i\cdot}^{(\alpha\beta)}\bar{R}_{i\cdot}^{(\alpha\gamma)},$ $(r_{2}^{(\alpha\beta)})^{2}$ $=\frac{1}{N_{\alpha}(N_{\beta}-\delta_{\alpha\beta})}\sum_{i=1}^{N_{\alpha}}\sum_{j=1}^{N_{\beta}}({R}_{ij}^{(\alpha\beta)})^{2},$
@@
Here $\bar{R}_{i\cdot}^{(\alpha\beta)}$ is the average rank per row of $\mathbf{R}^{(\alpha\beta)}$; $r_{0}^{(\alpha\beta)}$ is the mean of $\bar{R}_{i\cdot}^{(\alpha\beta)}$’s, which is also the average rank among $\mathbf{R}^{(\alpha\beta)}$; $(r_{2}^{(\alpha\beta)})^{2}$ and $(r_{1}^{(\alpha\beta\gamma)})^{2}$ are the second-order moments; $V_{2}^{(\alpha\beta)}$ and $V_{1}^{(\alpha\beta\gamma)}$ are two types of variance term corresponding to $r_{2}^{(\alpha\beta)}$ and $r_{1}^{(\alpha\beta\gamma)}$. We establish several useful properties of these quantities, valid for any symmetric matrix $\mathbf{R}$ with zero diagonal; these properties simplify the calculation of $\boldsymbol{\Sigma}_{v}$ and $\boldsymbol{\Sigma}_{c}$. Details are given in Supplementary C.
Let $\mathcal{T}_{x}^{(\alpha,\beta)}=n_{\alpha}(m_{\beta}-1-\delta_{\alpha\beta})(N_{\beta}-\delta_{\alpha\beta})V_{1}^{(\alpha\beta\beta)},$ $\mathcal{T}_{y}^{(\alpha,\beta)}=m_{\alpha}(n_{\beta}-1-\delta_{\alpha\beta})(N_{\alpha}-\delta_{\alpha\beta})V_{1}^{(\beta\alpha\alpha)},$ and $\mathcal{T}^{(\alpha,\beta)}=(N_{\beta}-\delta_{\alpha\beta})V_{1}^{(\alpha\beta\beta)}$. Theorem 3 gives the explicit expression of basic quantities that are used to construct $\boldsymbol{\Sigma}_{v}$ and $\boldsymbol{\Sigma}_{c}$ in our test statistics.
@@
Using the result in Theorem 3, we can directly obtain the BRISE-v statistic. While for calculating BRISE-c, we can utilize the following relations:
@@
The proof of Theorem 3 is given in Supplementary D. Particularly, the moments $\mathbb{E}(U_{x}^{(\alpha\beta)})$, $\mathbb{E}(U_{y}^{(\alpha\beta)})$, $\mathrm{Var}(U_{x}^{(\alpha\beta)}),\mathrm{Var}(U_{y}^{(\alpha\beta)})$ and $\mathrm{Cov}(U_{x}^{(\alpha\beta)},U_{y}^{(\alpha\beta)})$ will degenerate to the result in Theorem 1 of Zhou and Chen (2023) when $\alpha=\beta$ (i.e., two patterns are the same); RISE can also be viewed as the special case with a single pattern in our proposed framework.
@@ 3.2 Asymptotic Properties
Exact $p$-value computation via full permutation is feasible for small samples but becomes intractable as sample size grows. To enable scalable inference, we derive the asymptotic distributions of the test statistics $T_{v}$ and $T_{c}$ under the null. We use the following asymptotic notations: $a_{n}\prec b_{n}$ indicates $a_{n}$ is dominated by $b_{n}$; $a_{n}\asymp b_{n}$ denotes boundedness above and below; and $a_{n}\precsim b_{n}$ denotes boundedness above. We adopt a standard regime where $m_{\alpha},n_{\alpha}\to\infty$, with $\lim_{N\to\infty}N_{\alpha}/N\in(0,1)$ and $m_{\alpha}/N_{\alpha}\to p_{\alpha}\in(0,1)$ for all $1\leq\alpha\leq n_{P}$.
To introduce our main result, we further define the following notations. For each pattern pair $(\mathcal{Z}^{(\alpha)},\mathcal{Z}^{(\beta)})$, define the centered row sum $\tilde{R}_{i\cdot}^{(\alpha\beta)}=\bar{R}_{i\cdot}^{(\alpha\beta)}-r_{0}^{(\alpha\beta)}$. To capture the cross-pattern second-order dependence, we define
@@
In the standard limit regime, for all $(\alpha,\beta),(\beta,\gamma)\in\mathcal{I}$, under Conditions
$(1)\,$ $r_{1}^{(\alpha\beta\beta)}\prec r_{2}^{(\alpha\beta)};\qquad(2)\,\sum_{i=1}^{N_{\alpha}}(\sum_{j=1}^{N_{\beta}}(R_{ij}^{(\alpha\beta)})^{2})^{2}\lesssim N_{\alpha}^{2}N_{\beta}(r_{2}^{(\alpha\beta)})^{4};$
$(3)\,$ $\sum_{i=1}^{N_{\alpha}}\left|\tilde{R}_{i\cdot}^{(\alpha\beta)}\right|^{3}\prec(N_{\alpha}V_{1}^{(\alpha\beta\beta)})^{1.5};\qquad(4)\,\sum_{i=1}^{N_{\alpha}}\left|\tilde{R}_{i\cdot}^{(\alpha\beta)}\right|^{3}\prec N_{\alpha}r_{2}^{(\alpha\beta)}V_{1}^{(\alpha\beta\beta)};$
$(5)\,$ $\Big{|}\sum_{i=1}^{N_{\alpha}}\sum_{j=1}^{N_{\beta}}\sum_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{N_{\beta}}R_{ij}^{(\alpha\beta)}R_{il}^{(\alpha\beta)}\tilde{R}_{j\cdot}^{(\beta\gamma)}\tilde{R}_{l\cdot}^{(\beta\gamma)}\Big{|}\prec N_{\alpha}N_{\beta}^{2}(r_{2}^{(\alpha\beta)})^{2}V_{1}^{(\beta\gamma\gamma)};$
$(6)\,$ $\sum_{i=1}^{N_{\alpha}}\sum_{j=1}^{N_{\beta}}\sum_{\begin{subarray}{c}l=1\\ l\neq i,\,j\delta_{\alpha\beta}\end{subarray}}^{N_{\alpha}}\sum_{\begin{subarray}{c}s=1\\ s\neq j,\,i\delta_{\alpha\beta}\end{subarray}}^{N_{\beta}}R_{ij}^{(\alpha\beta)}R_{lj}^{(\alpha\beta)}R_{is}^{(\alpha\beta)}R_{ls}^{(\alpha\beta)}\prec N_{\alpha}^{2}N_{\beta}^{2}(r_{2}^{(\alpha\beta)})^{4}.$
and (C1): $\mathbf{E_{\alpha}}$ is positive definite, we have
@@
The proof of Theorem 4 is given in Supplementary E. The validity of Conditions (1)–(6) depends on the value of $\alpha$, $\beta$, and $\gamma$. When $\alpha=\beta=\gamma$, the conditions degenerate to the conditions given in Theorem 4 in Zhou and Chen (2023), where it was argued that the conditions are mild and plausibly met in real data settings. When $\alpha,\beta,\gamma$ are not all the same, Conditions (1), (2), (4), and (6) hold assuming that there are no nodes with a degree larger than $Ck$ for some constant $C$ in the pair-wise $k$-NNG defined in Section 2.2. To see this, let $G^{(\alpha\beta)}$ be the $k$-NNG under pattern pair $(\mathcal{Z}^{(\alpha)},\mathcal{Z}^{(\beta)})$, then we have $k=\max_{\alpha\beta,ij}R_{ij}^{(\alpha\beta)}$, and adopt the assumption that $k\prec N$ for the graph-induced rank. With a slight abuse of notation, we use $|\cdot|$ to denote the total weight of a graph. Usually we have $r_{0}^{(\alpha\beta)}\asymp k|G^{(\alpha\beta)}|/N_{\alpha}N_{\beta}$ and $(r_{2}^{(\alpha\beta)})^{2}\asymp k^{2}|G^{(\alpha\beta)}|/N_{\alpha}N_{\beta}$ where $|G^{(\alpha\beta)}|\asymp(N_{\alpha}+N_{\beta})k/2$. Thus, we have that Conditions (1), (2), (4), and (6) always hold as
@@
$\sum_{i=1}^{N_{\alpha}}{|\tilde{R}_{i\cdot}^{(\alpha\beta)}|^{3}}\leq\max_{i}|\tilde{R}_{i\cdot}^{(\alpha\beta)}|N_{\alpha}V_{1}^{(\alpha\beta\beta)}\lesssim\frac{k^{2}}{N_{\alpha}}N_{\alpha}V_{1}^{(\alpha\beta\beta)}\prec N_{\alpha}r_{2}^{(\alpha\beta)}V_{1}^{(\alpha\beta\beta)},$
@@
We further verify Conditions (3) and (5) by simulations. In our simulations, we calculate the ratio of the LHS to RHS of the inequality. Together with Condition (C1), simulation details and results are summarized in Supplementary N. The simulation results support that these conditions are easily satisfied.
Next, we discuss the consistency of BRISE. For the consistency under the standard limiting regime (see Theorem 5), Conditions (1)–(6) are not required. We consider $G^{(\alpha\beta)}$ constructed using Euclidean distance. When $k=O(1)$ and $N_{\alpha},N_{\beta}\to\infty$, each edge in $G^{(\alpha\beta)}$ exhibits asymptotically vanishing length, with connected nodes converging to arbitrarily close proximity in space (Henze and Penrose, 1999). This property arises because, in the limit, any point possesses sufficiently many neighbors within an arbitrarily small neighborhood to constitute its nearest neighbors. Under the choice of Euclidean distance, we have the following theorem.
@@
The proof of Theorem 5 is detailed in Supplementary F, following the methodologies of Schilling (1986) and Henze and Penrose (1999). This result may extend to alternative dissimilarity measures, provided they exhibit the same limiting behavior as described for $G^{(\alpha\beta)}$.
While Theorem 5 establishes the consistency of BRISE in the standard asymptotic settings, many modern applications, like genomics, neuroimaging, and text data, operate in the high-dimensional, low-sample-size (HDLSS) regime. In this context, classical asymptotic theory may fail due to the curse of dimensionality and degeneracy of distance metrics. The following result extends the consistency of BRISE to the HDLSS setting under mild moment and separation conditions.
@@
$\frac{1}{d_{\alpha\beta}}\mathbb{E}(\rho(X^{(\alpha)},\mathbb{E}X^{(\beta)})^{2})\to\sigma_{x,\alpha\beta}^{2},\frac{1}{d_{\alpha\beta}}\mathbb{E}(\rho(Y^{(\alpha)},\mathbb{E}Y^{(\beta)})^{2})\to\sigma_{y,\alpha\beta}^{2},\frac{1}{d_{\alpha\beta}}\mathbb{E}(\rho(X^{(\alpha)},\mathbb{E}Y^{(\beta)})^{2})\to\nu_{\alpha\beta}^{2},$
@@
    The largest eigenvalue of the covariance matrix $\boldsymbol{\Sigma}_{v}$ satisfies $\lambda_{+}[\boldsymbol{\Sigma}_{v}]=O(N^{2-b})$ for some $b>0$, $N\geq C_{b,\theta}$ for a constant $C_{b,\theta}>0$ only depending on $b,\theta$.
Assumptions (i)-(iii) of Theorem 6 are extended from Biswas and Ghosh (2014), with the proof of the theorem provided in Supplementary G. Here we only consider the overlapping Euclidean distance case of the dissimilarity measure given in (1). If using other measuring methods, our proposed tests may still be consistent under HDLSS as long as they have similar limiting constraints under suitable conditions.
@@ 4 Simulation Studies
We conduct simulation studies to evaluate the empirical type-I error and power of BRISE. We consider both BRISE-c and BRISE-v with Euclidean distance in (1) and $k=10$ for constructing the $k$-NNG, following the recommendation of Zhou and Chen (2023). We also examine the sensitivity of our methods to the choice of $k$ (see additional results reported in Supplementary O). The findings indicate that the performance of BRISE is stable with respect to $k$, as long as $k$ is not chosen too small. For both variants, $p$-values are obtained using their asymptotic null distributions.
We compare our methods with five baselines: MMD-Miss, RISE with $k$-NNG, the standard MMD test (Gretton et al., 2012), Ball Divergence (BD) (Pan et al., 2018), and Measure Transportation (MT) (Deb and Sen, 2021). Among these, only MMD-Miss can be directly applied to incomplete data; the others are implemented using complete-case analysis. The $p$-values of MMD-Miss, MMD, MT, and BD are obtained using $1000$ permutations, while RISE uses its asymptotic null distribution for $p$-value calculation. For RISE, we set $k=10$, and for the other baselines, we adopt their default hyperparameters. To further examine the advantage of the proposed pattern-wise permutation procedure in type-I error control, we also include a variant of BRISE-c that differs only in the permutation scheme, using standard instead of pattern-wise permutation, denoted as BRISE-c(sp). According to Theorem 1, this variant may fail when the marginal pattern probabilities differ between groups.
Data are generated as follows. For each observation, every source $Z^{(l)}\in\mathbb{R}^{d_{l}}$ is generated and observed independently with probability $p_{g,l}$ for $g\in{X,Y}$; observations with all sources missing are discarded. To reflect a variety of realistic scenarios, we simulate a wide range of null and alternative distributions in moderate to high-dimensional settings. These include multivariate Gaussian, log-normal, and $t_{5}$ distributions, covering both light-tailed and heavy-tailed behaviors. The alternative hypotheses encompass differences in location, scale, and their combinations.
We further define $d=\sum_{l=1}^{L}d_{l}$. For simplicity, we assume that all sources within the same group $g\in\{X,Y\}$ share a common sampling rate $p_{g}$, and that the $d_{l}$’s are identical across sources. Specifically, we consider the following source distributions:
@@
      Location: $\mathcal{N}_{d}(\boldsymbol{\mu},\Sigma_{X})$, where $\boldsymbol{\mu}=0.4\log{d}\cdot\boldsymbol{\mu}^{\prime}/\|\boldsymbol{\mu}^{\prime}\|_{2}$ and $\boldsymbol{\mu}^{\prime}\sim\mathcal{N}_{d}(0_{d},I_{d})$.
@@
      Scale: $\mathcal{N}_{d}(\mathbf{0}_{d},\sigma^{2}\Sigma_{X})$, with $\sigma=1+0.12\log{d}/\sqrt{{d}}$.
@@
      Location and scale: $\mathcal{N}_{d}(\boldsymbol{\mu},\Sigma_{Y})$, where $\boldsymbol{\mu}=0.1\log{d}\cdot\boldsymbol{\mu}^{\prime}/\|\boldsymbol{\mu}^{\prime}\|_{2}$, $\boldsymbol{\mu}^{\prime}\sim\mathcal{N}_{d}(0_{d},I_{d})$, and $\Sigma_{Y,ij}=0.3^{|i-j|}$.
@@
      Location: $\exp(\mathcal{N}_{d}(\boldsymbol{\mu},\Sigma_{X}))$, where $\mu_{j}=(-1)^{j}\cdot 2\log{d}/\sqrt{{d}}$ for $j=1,\dots,\lfloor 0.05{d}\rfloor$, and $\mu_{j}=0$ otherwise.
@@
      Scale: $\exp(\mathcal{N}_{d}(\mathbf{0}_{d},\sigma^{2}\Sigma_{X}))$, with $\sigma=1+0.15\log{d}/\sqrt{{d}}$.
@@
      Location and scale: $\exp(\mathcal{N}_{d}(\mu\mathbf{1}_{d},\sigma\Sigma_{X}))$, where $\mu=0.25\log{d}/\sqrt{{d}}$ and $\sigma=1+0.1\cdot(50/{d})^{0.25}$.
@@
    Multivariate Student-$t$. $(X^{(1)},\ldots,X^{(L)})\sim t_{5}(\mathbf{0}_{d},\Sigma_{X})$, where $\Sigma_{X,ij}=0.6^{|i-j|}$. $(Y^{(1)},\ldots,Y^{(L)})\sim F_{Y}(\cdot)$, where $F_{Y}(\cdot)$ is:
@@
      Location: $t_{5}(\boldsymbol{\mu},\Sigma_{X})$, where $\mu_{j}=(-1)^{j}\cdot 2\log{d}/\sqrt{{d}}$ for $j=1,\dots,\lfloor 0.05{d}\rfloor$, and $\mu_{j}=0$ otherwise.
@@
      Scale: $t_{5}(\mathbf{0}_{d},\Sigma_{Y})$, where $\Sigma_{Y,ij}=0.75\cdot 0.3^{|i-j|}$.

### v2

@@ 4.1 Finite-sample Properties
In this section, we provide explicit formulas for $\boldsymbol{\Sigma}_{v}$ and $\boldsymbol{\Sigma}_{c}$ under the null hypothesis, which are crucial for constructing the proposed test statistics.
We first introduce notation related to the rank matrix. Let $\delta_{\alpha\beta}=\mathbbm{1}(\alpha=\beta)$ be the Kronecker delta with the indicator function $\mathbbm{1}(\cdot)$. We further define:
$\widebar{R}_{i\cdot}^{(\alpha\beta)}$ $=\frac{1}{N_{\beta}-\delta_{\alpha\beta}}\sum_{j=1}^{N_{\beta}}R_{ij}^{(\alpha\beta)},$ $r_{0}^{(\alpha\beta)}$ $=\frac{1}{N_{\alpha}}\sum_{i=1}^{N_{\alpha}}\widebar{R}_{i\cdot}^{(\alpha\beta)},$
$(r_{1}^{(\alpha\beta\gamma)})^{2}$ $=\frac{1}{N_{\alpha}}\sum_{i=1}^{N_{\alpha}}\widebar{R}_{i\cdot}^{(\alpha\beta)}\widebar{R}_{i\cdot}^{(\alpha\gamma)},$ $(r_{2}^{(\alpha\beta)})^{2}$ $=\frac{1}{N_{\alpha}(N_{\beta}-\delta_{\alpha\beta})}\sum_{i=1}^{N_{\alpha}}\sum_{j=1}^{N_{\beta}}({R}_{ij}^{(\alpha\beta)})^{2},$
@@
Here $\widebar{R}_{i\cdot}^{(\alpha\beta)}$ denotes the average of the ranks in the $i$th row of $\mathbf{R}^{(\alpha\beta)}$; $r_{0}^{(\alpha\beta)}$ is the mean of $\widebar{R}_{i\cdot}^{(\alpha\beta)}$’s, which is also the average rank among $\mathbf{R}^{(\alpha\beta)}$; $(r_{2}^{(\alpha\beta)})^{2}$ and $(r_{1}^{(\alpha\beta\gamma)})^{2}$ are the second-order moments; $V_{2}^{(\alpha\beta)}$ and $V_{1}^{(\alpha\beta\gamma)}$ are two types of variance term corresponding to $r_{2}^{(\alpha\beta)}$ and $r_{1}^{(\alpha\beta\gamma)}$. We establish several useful properties of these quantities for any symmetric matrix $\mathbf{R}$ with zero diagonal; these properties simplify the calculation of $\boldsymbol{\Sigma}_{v}$ and $\boldsymbol{\Sigma}_{c}$. Details are given in Supplementary C.
Let $\mathcal{T}_{x}^{(\alpha,\beta)}=n_{\alpha}(m_{\beta}-1-\delta_{\alpha\beta})(N_{\beta}-\delta_{\alpha\beta})V_{1}^{(\alpha\beta\beta)},$ $\mathcal{T}_{y}^{(\alpha,\beta)}=m_{\alpha}(n_{\beta}-1-\delta_{\alpha\beta})(N_{\alpha}-\delta_{\alpha\beta})V_{1}^{(\beta\alpha\alpha)},$ and $\mathcal{T}^{(\alpha,\beta)}=(N_{\beta}-\delta_{\alpha\beta})V_{1}^{(\alpha\beta\beta)}$. Theorem 3 provides explicit expressions for the quantities used to construct $\boldsymbol{\Sigma}_{v}$ and $\boldsymbol{\Sigma}_{c}$.
@@
Using the result in Theorem 3, we can directly calculate the BRISE-v statistic. While for calculating BRISE-c, we further utilize the following relations:
@@
In particular, when $\alpha=\beta$, these moments reduce to those in Theorem 1 of Zhou and Chen (2023), corresponding to the fully observed setting. Thus, RISE can be viewed as a special case of the proposed framework with fully observed data.
@@ 4.2 Asymptotic Properties
Exact $p$-value computation via full pattern-wise permutation is feasible for small samples but becomes computationally intractable as sample size grows. To enable scalable inference, we derive the asymptotic distributions of the test statistics $T_{v}$ and $T_{c}$ under the null. We use the following asymptotic notations: $a_{n}\prec b_{n}$ indicates $a_{n}$ is dominated by $b_{n}$; $a_{n}\asymp b_{n}$ denotes boundedness above and below; and $a_{n}\precsim b_{n}$ denotes boundedness above. We adopt a standard asymptotic regime where $m_{\alpha},n_{\alpha}\to\infty$, with $\lim_{N\to\infty}N_{\alpha}/N\in(0,1)$ and $m_{\alpha}/N_{\alpha}\to p_{\alpha}\in(0,1)$ for all $1\leq\alpha\leq n_{P}$.
To introduce our main result, we further define the following notations. For each pattern pair $(\mathcal{Z}^{(\alpha)},\mathcal{Z}^{(\beta)})$, define the centered row sum $\widetilde{R}_{i\cdot}^{(\alpha\beta)}=\widebar{R}_{i\cdot}^{(\alpha\beta)}-r_{0}^{(\alpha\beta)}$. To capture the cross-pattern second-order dependence, we define
@@
[…]
@@
The validity of Conditions (1)–(6) depends on the indices $\alpha$, $\beta$, and $\gamma$. When $\alpha=\beta=\gamma$, these conditions reduce to those in Theorem 4 of Zhou and Chen (2023), which were shown to be mild and plausible in practice. When $\alpha,\beta,\gamma$ are not all the same, Conditions (1), (2), (4), and (6) hold assuming that there are no nodes with a degree larger than $Ck$ for some constant $C$ in the pair-wise $k$-NNG defined in Section 3. To see this, let $G^{(\alpha\beta)}$ be the $k$-NNG given a pair of missing patterns $(\mathcal{Z}^{(\alpha)},\mathcal{Z}^{(\beta)})$, then we have $k=\max_{\alpha\beta,ij}R_{ij}^{(\alpha\beta)}$, and adopt the assumption that $k\prec N$ for the graph-induced rank. With a slight abuse of notation, we use $|\cdot|$ to denote the total amount of edges in a graph. Usually we have $r_{0}^{(\alpha\beta)}\asymp k|G^{(\alpha\beta)}|/N_{\alpha}N_{\beta}$ and $(r_{2}^{(\alpha\beta)})^{2}\asymp k^{2}|G^{(\alpha\beta)}|/N_{\alpha}N_{\beta}$ where $|G^{(\alpha\beta)}|\asymp(N_{\alpha}+N_{\beta})k/2$. Thus, we have that Conditions (1), (2), (4), and (6) always hold as
@@
$\sum_{i=1}^{N_{\alpha}}{|\widetilde{R}_{i\cdot}^{(\alpha\beta)}|^{3}}\leq\max_{i}|\widetilde{R}_{i\cdot}^{(\alpha\beta)}|N_{\alpha}V_{1}^{(\alpha\beta\beta)}\lesssim\frac{k^{2}}{N_{\alpha}}N_{\alpha}V_{1}^{(\alpha\beta\beta)}\prec N_{\alpha}r_{2}^{(\alpha\beta)}V_{1}^{(\alpha\beta\beta)},$
@@
We further verify Conditions (3) and (5) via simulation by evaluating the ratio of the left-hand side to the right-hand side of each inequality. Together with Condition (C1), simulation details and results are summarized in Supplementary O. The simulation results support that these conditions are easily satisfied.
We next establish the consistency of BRISE. For the consistency under the standard limiting regime (see Theorem 5), Conditions (1) - (6) are not required. We consider $G^{(\alpha\beta)}$ constructed using Euclidean distance. When $k=O(1)$ and $N_{\alpha},N_{\beta}\to\infty$, each edge in $G^{(\alpha\beta)}$ exhibits asymptotically vanishing length, with connected nodes converging to arbitrarily close proximity in space (Henze and Penrose, 1999). This property arises because, in the limit, any point possesses sufficiently many neighbors within an arbitrarily small neighborhood to constitute its nearest neighbors. Under the choice of Euclidean distance, we have the following theorem.
@@
The proof of Theorem 5 is detailed in Supplementary F, following the methodologies of Schilling (1986) and Henze and Penrose (1999). This result may extend to alternative dissimilarity measures, provided they exhibit the same limiting behavior as $G^{(\alpha\beta)}$.
While Theorem 5 establishes the consistency of BRISE in the standard asymptotic settings, many modern applications, like genomics, neuroimaging, and text data, operate in the high-dimensional, low-sample-size (HDLSS) regime. In this context, classical asymptotic theory may fail due to the curse of dimensionality and the degeneracy of distance metrics. The following result extends the consistency of BRISE to the HDLSS setting under mild moment and separation conditions.
@@
$\frac{1}{d_{\alpha\beta}}\rho(\mathbb{E}(\rho((X^{(\alpha)},\mathbb{E}X^{(\beta)})^{2})\to\sigma_{x,\alpha\beta}^{2},\qquad\frac{1}{d_{\alpha\beta}}\mathbb{E}(\rho(Y^{(\alpha)},\mathbb{E}Y^{(\beta)})^{2})\to\sigma_{y,\alpha\beta}^{2},$
$\frac{1}{d_{\alpha\beta}}\mathbb{E}(\rho(\mathbb{E}X^{(\alpha)},\mathbb{E}Y^{(\beta)})^{2})\to\nu_{\alpha\beta}^{2},$
@@
    The largest eigenvalue of the covariance matrix $\boldsymbol{\Sigma}_{v}$, denoted $\lambda_{+}[\boldsymbol{\Sigma}_{v}]$, is bounded by $\lambda_{+}[\boldsymbol{\Sigma}_{v}]\leq cN^{2-b}$ for constants $c,b>0$, and $N\geq C_{b,\theta}$ where $C_{b,\theta}>0$ is a constant only depending on $b,\theta$.
Assumptions (i) - (iii) of Theorem 6 are extended from Biswas and Ghosh (2014). Here we focus on the overlapping Euclidean distance case in (1). The proposed tests may remain consistent under other dissimilarity measures, provided analogous limiting conditions hold.
@@ 5 Simulation Studies
We conduct simulation studies to evaluate the validity of BPET framework, and empirical type-I error and power of BRISE. We consider both BRISE-c and BRISE-v with Euclidean distance in (1) and $k=10$ for constructing the $k$-NNG, following the recommendation of Zhou and Chen (2023). We also examine the sensitivity of our methods to the choice of $k$ (see additional results reported in Supplementary P). The findings indicate that the performance of BRISE is stable with respect to $k$, as long as $k$ is not chosen too small. For both variants, $p$-values are obtained using their asymptotic null distributions.
We compare our methods with five baselines: MMD-Miss, RISE with $10$-NNG, the standard MMD test (Gretton et al., 2012), Ball Divergence (BD) (Pan et al., 2018), and Measure Transportation (MT) (Deb and Sen, 2021), with their default hyperparameters. Among these, only MMD-Miss can be directly applied to incomplete data; the others are implemented via complete-case analysis. The $p$-values of MMD-Miss, MMD, MT, and BD are obtained using $1000$ permutations, while RISE uses its asymptotic null distribution for $p$-value calculation.
To further assess whether imputation can serve as a viable alternative under block-wise missingness, we also consider an imputation-based strategy in low-dimensional settings: we first apply MICE (Van Buuren and Groothuis-Oudshoorn, 2011) to impute the missing data, and then apply RISE, MMD, BD, and MT, respectively, to the imputed datasets. This comparison is restricted to $d\in\{10,20,50\}$ due to the computational cost and instability of multivariate imputation in higher dimensions.
To examine the advantage of the proposed pattern-wise permutation procedure in type-I error control, we further include a variant of BRISE-c that differs only in the permutation scheme, using standard instead of pattern-wise permutation, denoted as BRISE-c(sp). According to Theorem 1, this variant may fail when the marginal pattern probabilities differ between groups. We emphasize that most competing methods are not designed for block-wise missingness and therefore require either complete-case analysis or model-based imputation. Our goal is not to advocate a particular imputation model, but to provide a test that operates directly on observed data without imputation.
To simulate block-wise missing data, for each observation, each source/modality $Z^{(l)}\in\mathbb{R}^{d_{l}}$ is generated and observed independently with probability $p_{G}$ for $G\in\{X,Y\}$; observations with all sources missing are discarded. In total, we have $d=\sum_{l=1}^{L}d_{l}$ variables. To reflect a variety of realistic scenarios, we simulate a wide range of null and alternative distributions in moderate- to high-dimensional settings, including multivariate Gaussian, log-normal, and $t_{5}$ distributions, covering both light-tailed and heavy-tailed behaviors. The alternatives include location, scale, and mixed differences. Specifically, we consider the following source/modality distributions:
@@
      Location: $\mathcal{N}_{d}(\boldsymbol{\mu},\Sigma_{X})$, where $\boldsymbol{\mu}=0.4\log{d}\cdot\boldsymbol{\mu}^{\prime}/\|\boldsymbol{\mu}^{\prime}\|_{2}$ and $\boldsymbol{\mu}^{\prime}\sim\mathcal{N}_{d}(0_{d},I_{d})$;
@@
      Scale: $\mathcal{N}_{d}(\mathbf{0}_{d},\sigma^{2}\Sigma_{X})$, with $\sigma=1+0.12\log{d}/\sqrt{{d}}$;
@@
      Mixed: $\mathcal{N}_{d}(\boldsymbol{\mu},\Sigma_{Y})$, where $\boldsymbol{\mu}=0.1\log{d}\cdot\boldsymbol{\mu}^{\prime}/\|\boldsymbol{\mu}^{\prime}\|_{2}$, $\boldsymbol{\mu}^{\prime}\sim\mathcal{N}_{d}(0_{d},I_{d})$, and $\Sigma_{Y,ij}=0.3^{|i-j|}$.
@@
      Location: $\exp(\mathcal{N}_{d}(\boldsymbol{\mu},\Sigma_{X}))$, where $\mu_{j}=(-1)^{j}\cdot 2\log{d}/\sqrt{{d}}$ for $j=1,\dots,\lfloor 0.05{d}\rfloor$, and $\mu_{j}=0$ otherwise;
@@
      Scale: $\exp(\mathcal{N}_{d}(\mathbf{0}_{d},\sigma^{2}\Sigma_{X}))$, with $\sigma=1+0.15\log{d}/\sqrt{{d}}$;
@@
      Mixed: $\exp(\mathcal{N}_{d}(\mu\mathbf{1}_{d},\sigma\Sigma_{X}))$, where $\mu=0.25\log{d}/\sqrt{{d}}$ and $\sigma=1+0.1\cdot(50/{d})^{0.25}$.
@@
    Multivariate $t_{5}$. $(X^{(1)},\ldots,X^{(L)})\sim t_{5}(\mathbf{0}_{d},\Sigma_{X})$, where $\Sigma_{X,ij}=0.6^{|i-j|}$. $(Y^{(1)},\ldots,Y^{(L)})\sim F_{Y}(\cdot)$, where $F_{Y}(\cdot)$ is:
@@
      Location: $t_{5}(\boldsymbol{\mu},\Sigma_{X})$, where $\mu_{j}=(-1)^{j}\cdot 2\log{d}/\sqrt{{d}}$ for $j=1,\dots,\lfloor 0.05{d}\rfloor$, and $\mu_{j}=0$ otherwise;
@@
      Scale: $t_{5}(\mathbf{0}_{d},\Sigma_{Y})$, where $\Sigma_{Y,ij}=0.75\cdot 0.3^{|i-j|}$;